OpenStudy (aye_my_peeps):

evaluate: 3-7 times 5+2 times 4. a. -40 b. -24 c. -12 d. 28

OpenStudy (aye_my_peeps):

@ScarlettFarra2000

563blackghost (563blackghost):

So we have the equation \(\Huge{3-7 \times 5+2 \times 4}\) Well we would use PEMDAS what part of the equation would we do first?

OpenStudy (aye_my_peeps):

multiply

OpenStudy (aye_my_peeps):

i got -40 am i right??

OpenStudy (aye_my_peeps):

hello??

563blackghost (563blackghost):

sorry my internet disconnected :( Yes we would multiply first.... \(\LARGE{7 \times 5 =?}\) \(\LARGE{2 \times 4=?}\) What would these two be?

563blackghost (563blackghost):

@AYE_MY_PEEPS

OpenStudy (aye_my_peeps):

35 and 8

563blackghost (563blackghost):

Correct so we now input we look at our equation again.... \(\Huge{3-35+8}\) Now to solve this we would first subtract since we solve from left to right...so what is \(\LARGE{3-35}\) ?

OpenStudy (aye_my_peeps):

-32

563blackghost (563blackghost):

Correct so we are left with \(\Huge{-32+8}\) So what is our answer?

OpenStudy (aye_my_peeps):

-24

563blackghost (563blackghost):

Correct ^^

OpenStudy (aye_my_peeps):

thnx :)

563blackghost (563blackghost):

Np :D